OpenStudy (anonymous):

solve -5√x-9-7=23? (radical equation). answers: x = 3 x = 45 x = –6 None of the above

OpenStudy (valpey):

What you have written looks like the square root function would have to return a negative number, which isn't possible by the definition of the square root function, so None of the above. That said, I'm not sure we can read the problem completely. Did it look like this: \[-5\sqrt{x-9}-7=23\] or \[-5\sqrt{x}-9-7=23\] or \[-5\sqrt{x-9-7}=23\] or\[\sqrt[5]{x-9}-7=23\]

OpenStudy (anonymous):

the first one.

OpenStudy (valpey):

And you are sure it wasn't: \[5\sqrt{x-9}-7=23\]

OpenStudy (anonymous):

Okay I looked at it again, and yes it is that one

OpenStudy (anonymous):

the 5 is definatly not an exponent tho.

OpenStudy (valpey):

Okay, good, that would be tricky. (the takeaway from that distinction is that you can't take the square root of something and get a negative number). But here is how you solve this: \[5\sqrt{x-9}-7=23\]add seven to both sides: \[5\sqrt{x-9}=30\]divide both sides by 5: \[\sqrt{x-9}=6\]square both sides: \[x-9=36\] add 9 to both sides: \[x=45\]

OpenStudy (anonymous):

thank you very much, that was helpful.

OpenStudy (valpey):

You bet. Good luck with other questions. Just take it one step at a time (...what can I do to both sides of the equation to make it simpler?)

OpenStudy (anonymous):

okay, :) again ty.

OpenStudy (radar):

\[-5\sqrt{x-9}-7=23\]\[-5\sqrt{x-9}=30\]\[\sqrt{x-9}=-6\]\[x-9=36\]\[x=45\] The root of a number may be negative or positive or some cases both \[\sqrt[3]-8{}=-2\]

OpenStudy (anonymous):

okay ty2u2.

OpenStudy (radar):

\[\sqrt{4}=\pm2\]

OpenStudy (anonymous):

ah-ha. would you like to help me on a another problem? it is my last one for the night?

OpenStudy (radar):

I will try.

OpenStudy (anonymous):

okay ty.

OpenStudy (radar):

Even if my tractor is a Ford lol

OpenStudy (anonymous):